# PersonaFlow: Boosting Research Ideation with LLM-Simulated Expert Personas

### v1

@@ 3.2. System Interaction Design
Informed by the findings from our formative studies, we designed multiple features of the PersonaFlow system to address each design requirement. The overall system workflow follows an interactive ideation process, as described in Figure 1. PersonaFlow offers five major features:
@@
  • Persona Node: node representing the profile of each expert persona, with a customizer interface allowing users to edit the traits of each persona initially populated by the system.
@@
We provide more detailed descriptions for each feature in Section 3.2.2.
3.2.1. Example Use Case Scenario — a System Walkthrough
Imagine Linh, an HCI researcher specializing in Health Informatics, who recently grew interested in exploring the potential application of LLMs in their familiar domain. Since they are from the background of HCI researchers specializing in Health Informatics, Linh found it challenging to identify relevant literature on LLMs and synthesize it into meaningful research questions due to the interdisciplinary nature of the topic and the vast amount of information available.
Linh started by creating a new Research Question (RQ) Node in the mind map canvas to guide her exploration of LLMs in Health Informatics. With their initial idea of “fine-tuning Large Language Models to promote preventive health behavior,” Linh wrote down the idea in the initial RQ node and clicked the action button on the node to generate follow-up nodes. The system then generated three Persona Nodes based on the initial idea provided by Linh, which were presented on the canvas and connected to the initial RQ Node.
After examining the details of the three expert personas, Linh found one of the personas “Behavior Psychology Researcher” particularly interesting. Curious about the details of the persona, Linh clicked on the node to reveal an in-depth persona profile containing different traits describing information about the role and background of the expert persona. After reviewing and editing the persona profile (adding “building AI-based tools” to the “Experience” field), Linh decided to use the “Behavior Psychology Researcher” persona to dive deeper into relevant literature by generating more subsequent nodes. Three new literature nodes were revealed each containing a preview of a collection of papers suggested by the expert persona based on a specific topic. Interested in the topic displayed on one of the Literature Nodes “Promoting Health Behavior Change With AI,” Linh decided to expand this node by clicking on it, revealing a detailed panel that allowed them to review each paper in detail, and make changes to the paper list by adding or removing entries as necessary to align with her specific research objectives.
After editing the list of papers, Linh proceeded to generate more follow-up nodes from the Literature Node. The system then generated three critique nodes, which provided feedback on Linh’s initial idea based on the persona-suggested literature. Linh was intrigued by one of the critiques focused on “Theoretical Framework,” hinting that the inclusion of some specific psychology theories and concepts such as “Transtheoretical Model of Change (TTM)” or “Information-Motivation-Behavioral Skills (IMB)” could benefit the design and evaluation of the intervention in preventive health behavior.
By further requesting follow-up nodes on the Critique Node, Linh found the three newly generated RQ Nodes to be a revised version of their initial idea. Linh chose one of the RQ Nodes displaying “How can fine-tuning Large Language Models based on the Transtheoretical Model of Change promote preventive health behavior?” and clicked on it to reveal a Research Outline panel on the right side of their screen. The panel first showcased a literature review summarizing the key findings from past works in bullet points, and requested Linh to further define a more detailed research scenario. Inspired by the literature review, Linh wrote down a detailed research scenario “Fine-tuning LLMs to deliver personalized stage-targeted health behavior change messages.” Based on their input, the system further generated a research outline table with sections such as motivation, objectives, methodology, expected outcomes, and potential limitations of a proposed study. Linh then reviewed the generated research outline table and made several detailed adjustments and edits based on their own ideas through the editing feature provided by the research outline panel.
Figure 2. Persona Node and Customizer: The Persona Node interface provides an overview of each expert persona’s profile information including name and traits. Upon clicking the node, a Persona Customizer panel shows up on the right side of the screen displaying the detailed traits and allowing the user to customize them.
@@ 3.2.2. System Features
Persona Node [D2]: expert knowledge and specialization in human-AI collaborative ideation. Existing research (Ali et al., 2019; Dastmalchi et al., 2021) about Transactive Memory Systems has suggested the importance of clear awareness of specialization in a team ideation setting. Thus, we introduce the design of persona nodes to represent domain-specific expertise, allowing users to interact with expert personas with a clear awareness of their specializations and knowledge background. By utilizing expert personas, we aim to stimulate LLM’s knowledge and assist in bridging the research efforts of researchers from different backgrounds. This design intends to mitigate users’ cognitive load, and improve the quality of feedback obtained from each dedicated persona. The persona node contains information about the detailed profile of each expert persona described in traits, as shown in Figure 2. When a user clicks on the persona node, the system displays a persona customizer panel on the right side where they can review and edit the details of the traits belonging to each persona. We took inspiration from Benharrak et al. (2024)’s work when designing the schema of persona traits, allowing users to modify both the role and detailed expertise of each expert persona. We also provide a field for users to specify any instructions they might have for the persona during the ideation process. The system synthesizes each persona’s initial profile based on the existing contextual information during the ideation. The detailed method we used to generate the persona profiles is presented in section 3.3.1. Users can modify persona profiles at any time during exploration by adding, deleting, or editing traits based on their needs and preferences. We do not automatically update or overwrite cascading nodes when changes are made to personas, but users can manually regenerate additional nodes to reflect the updated persona traits.
Literature Node [D3]: supporting rapid literature discovery in various domains. Identifying literature groundings and research gaps is an essential step during the ideation stage of the research lifecycle (Foster, 2004; Palmer et al., 2009). To assist users in efficiently navigating the vast amount of relevant literature, especially in unfamiliar domains, we introduced the design of Literature Node, as shown in Figure 3. Each of the Literature Nodes presents a specific research topic, with a collection of scholarly publications proposed by the preceding expert persona. The collection of papers is retrieved from an online publication database using the pipeline described in Section 3.3.2. When a user clicks on the node, a right-side panel is revealed with detailed information about each publication. The user will be able to delete irrelevant papers or add additional papers using a search engine-like interface.
@@
Critique Node [D2]: gathering feedback from expert personas. Based on a Persona Node and Literature Node, the user will be able to generate a follow-up Critique Node, as shown in component 4 of Figure 1. The Critique Node contains a user-editable critique of the initial research idea based on the given Literature Node and impersonation of the expert persona from the Persona Node. The critique nodes are further used to illustrate the process of how peer feedback is given from the perspective of a persona other than the researcher using the system. We use the design of the three types of nodes to model the process of peer discussions and knowledge solicitation with researchers or experts from different disciplines and domains.
@@
RQ Node and Research Outline Panel [D1]: facilitating iterative research ideation and refinement through expert feedback. Figure 4 demonstrates the RQ Node and Research Outline Panel feature. This feature allows users to iteratively improve their research ideas by generating and refining research questions (RQs) and their corresponding outlines. The design decision of making the system RQ-driven is due to the crucial role of RQ in research development and its nature of being iteratively refined during research ideation (Elio et al., 2011; Foster, 2004). When a user selects an RQ Node, the system automatically generates a Research Outline panel, which provides a structured overview of a potential research project derived from the RQ. The Research Outline Panel contains five sections: research question, literature review, research scenario, potential research outline, and expected outcomes. The literature review section provides a summary of relevant literature, prompting the user to think critically about the existing research landscape. The system then encourages the user to specify a detailed research scenario, with several system-generated suggestions. Based on this input, the system dynamically generates a comprehensive research outline with key sections such as motivation, objectives, methodology, expected outcomes, and potential limitations. This interactive process allows users to iteratively refine their research outline by reviewing and incorporating system-generated suggestions and insights.
@@ 3.3. PersonaFlow System Backend and Implementation
In this section, we detail the implementation details of the PersonaFlow  system backend222system source code will be released upon acceptance. The backend is designed to provide on-demand results by generating and synthesizing expert personas using LLM, identifying relevant literature based on research questions and personas, providing critiques by impersonating the personas, and generating research outlines. The PersonaFlow system is developed as a web application in Typescript, utilizing ReactJS and TailwindCSS for the frontend. The interactive flow editor is implemented using React Flow [^3] https://github.com/wbkd/react-flow/. For the application backend, Python with FastAPI [^4] https://github.com/tiangolo/fastapi/ is employed as the RESTful API server framework. Our LLM engine of choice is GPT-4o, and we present our detailed prompts used in Section A.4.
3.3.1. Persona synthesis
The system generates different personas to provide relevant literature and critique, based on the research context. The personas can be generated under 3 different scenarios: 1) Initial User RQ: When an RQ is initially provided by the user, the system uses the RQ to prompt our LLM to generate three relevant personas that simulate real-world individuals to augment ideation and brainstorming; 2) Literature Node-based Persona: When a researcher is recommended a literature node relevant to their RQ, they can generate personas using the literature node to ensure the generation of relevant personas. The persona is generated by utilizing the abstracts from the previous literature, which are processed using a scientific sentence classification pipeline built with a fine-tuned BERT model (Cohan et al., 2019; Beltagy et al., 2019). This pipeline extracts the background, methodology, and conclusion of the research. Finally, it summarizes the literature node to provide context for generating personas; 3) Author Profiling-based Persona Generation: To leverage the expertise of scholarly authors whose papers were included in the literature node, we developed an author profiling-based persona generation. Using the Semantic Scholar API (Kinney et al., 2023), we extracted papers from authors whose work was being recommended in the literature node. We identified the top three papers for each first author (since first authors are typically the primary contributors and experts in the research area the paper addresses), based on cosine similarity555with a threshold score of .65 using FAISS (Douze et al., 2024). These papers were summarized and used as context to generate personas aiming to capture the expertise and background of the authors.
@@ 3.3.1. Persona synthesis
3.3.2. Persona-driven Paper Retrieval
We designed and implemented a persona-driven scholarly paper retrieval module to assist users in identifying the relevant papers with higher coverage. However, using a research question directly as a search query to retrieve papers from online databases often leads to very few or irrelevant results. The PersonaFlow system conducts two steps when retrieving relevant papers: 1) query decomposition to yield high coverage of papers under different sub-topics; 2) semantic-based re-ranking to reduce the number of papers and only present the topic relevant ones to the user based on user’s prior research question. During the query decomposition step, we prompt the LLM to first break down the user’s initial research question into primary queries that capture various aspects of the main topic. Then, each of the primary queries will be decomposed into several secondary sub-queries to optimize the search results using the Semantic Scholar search API (Kinney et al., 2023). A similar approach has also been found to be used in existing work for optimizing literature discovery results (Zheng et al., 2024). For instance, the research question “How to address the lack of engagement of users using online art platforms by simulating different AI personas?” can be broken down into primary queries including “engagement in online art appreciation” and “multi-persona simulation using AI.” Subsequently, each primary query will be further decomposed into secondary queries that are shorter keywords. For example, the primary query “engagement in online art appreciation” will be decomposed into “user engagement,” “art appreciation” and “AI persona.” After the search results for each primary and secondary query are gathered, they are aggregated and re-ranked using a dense retrieval model (Chen et al., 2024) that aligns the retrieved papers with the user’s original research question. We take the top-k most relevant papers based on semantic similarity and present them to the user. During the later user studies, we present the top 10 most relevant papers to the user.
@@ 4. User Studies Evaluating Co-Creation with PersonaFlow
  1. RQ1: How does interacting with multiple LLM-simulated expert personas affect users’ perception of PersonaFlow-generated feedback, as well as their overall user experience? (Perception)
  2. RQ2: How do users customize the LLM-recommended expert personas, and what is the relationship between PersonaFlow feature use and users’ perceived overall user experience? (Behavior)
  3. RQ3: What are the different scenarios and use cases in which users employ PersonaFlow for scientific ideation during open-ended exploration? (Sample Use Cases)
@@ 4.1. User Study Design
We conducted two user studies to evaluate the PersonaFlow system. All studies were completed remotely online over Zoom, where participants were asked to share their screens. Study procedures were approved by the IRB of the researchers’ institution. Participants were compensated at \$20 per hour for all user studies.
4.1.1. Study 1: task-based evaluation with single- and multi-persona.
@@
We designed a within-subject user study to understand the impact of LLM-simulated personas on users’ research ideation process and how they interact with these personas. In order to examine our system’s utility in supporting interdisciplinary research ideation, we recruited 21 participants from 8 different research institutions and different backgrounds (e.g., HCI, Bio-informatics, Machine Learning, etc.) with prior experience in conducting interdisciplinary research. Detailed information about participants’ backgrounds can be found in Section A.1.
@@
Specifically, we focus on studying users’ engagement with different AI expert personas through the experiment design with two different conditions: research ideation with single-persona and multi-persona. The study was conducted over two sessions, each representing one of the two conditions. Under the single-persona condition, the user was asked to interact with a single AI expert persona and complete a single iteration of research ideation. We instructed the user to pick one AI persona out of the first batch of LLM synthesized personas, gather literature, and receive critiques until the first set of revised research questions is reached. Subsequently, the user was asked to experience the multi-persona condition by building on the outcomes from the previous iteration and continue exploring by generating and involving multiple AI personas, until they reached the second set of research questions. After each of the two sessions, the user was provided with a post-session survey form for quantitative evaluation of their experience pertaining to that session. By the end of the study, we conducted an exit interview to gather qualitative feedback on their experience and rationalization of their interactions when using the system.
4.1.2. Study 2: open-ended free exploration.
@@
After the formal study session, we invited participants to an optional 30-minute free exploration session where they could freely use the system’s features with minimal guidance and no procedural constraints. Participants were free to build on their existing flow from the previous session or create new flows using another topic. In total, 10 participants enrolled in the open-ended exploratory study. This study session aims to provide a deeper understanding of how users would use the system in a real-world exploration setting, as well as the benefits and shortcomings of the PersonaFlow system.

### v2

@@ 3.2. System Interaction Design
Informed by the findings from our formative studies, we designed multiple features of the PersonaFlow system to address each design requirement. PersonaFlow aims to support exploration of interdisciplinary research topics and ideas by involving diverse perspectives through LLM-simulated domain experts. The system is designed to benefit research exploration of interdisciplinary researchers, especially those planning to explore unfamiliar domains. The overall system workflow follows an interactive ideation process, as described in Figure 1. PersonaFlow offers five major features:
@@
  • Persona Node: node representing the profile of each simulated expert, with a customizer interface allowing users to edit the characteristics of each expert initially populated by the system.
@@
We provide more detailed descriptions for each feature in Section 3.2.1. The system does not require prior research expertise, but users need an initial idea to begin exploration. The system is primarily designed to support exploration from a rough idea and provide diverse perspectives, ultimately helping the user narrow down to a more specific concept.
Figure 2. Persona Node and Customizer: The Persona Node interface provides an overview of each expert persona’s profile information including name and characteristics. Upon clicking the node, a Persona Customizer panel shows up on the right side of the screen displaying the detailed characteristics and allowing the user to customize them.
[…]
Three components: Persona Node ‘NLP Algorithm Specialist’, Persona Node ‘Information Organization Expert’ and Persona Customizer side panel for ‘NLP Algorithm Specialist’. Upon clicking Persona Node ‘NLP Algorithm Specialist’, the Persona Customizer side panel shows detailed Persona information which contains traits such as ‘Role’, ‘Goal’, ‘Domain’ and ‘Expertise’, allowing the user to customize based on user preference.
@@ 3.2.1. System Features
Persona Node [D2]: expert knowledge and specialization in human-AI collaborative ideation. Existing research (Ali et al., 2019; Dastmalchi et al., 2021) about Transactive Memory Systems has suggested the importance of clear awareness of specialization in a team ideation setting. Thus, we introduce the design of persona nodes to represent domain-specific expertise, allowing users to interact with LLM-simulated experts with a clear awareness of their specializations and knowledge background. By utilizing expert personas, we aim to stimulate the LLM’s knowledge and assist in bridging the research efforts of researchers from different backgrounds. This design intends to mitigate users’ cognitive load, and improve the quality of feedback obtained from the perspective of each dedicated expert. The persona node contains information about the detailed profile of each expert persona described in characteristics, as shown in Figure 2. When a user clicks on the persona node, the system displays a persona customizer panel on the right side where they can review and edit the details of the characteristics belonging to each expert. We took inspiration from Benharrak et al. (2024)’s work when designing the schema of expert traits, allowing users to modify both the role and detailed expertise of each expert. We also provide a field for users to specify any instructions they might have for the persona during the ideation process. The system synthesizes each persona’s initial profile based on the existing contextual information during the ideation. The detailed method we used to generate the persona profiles is presented in section 3.3.1. Users can modify persona profiles at any time during exploration by adding, deleting, or editing characteristics based on their needs and preferences. We do not automatically update or overwrite cascading nodes when changes are made to personas, but users can manually regenerate additional nodes to reflect the updated persona characteristics.
Literature Node [D3]: rapid literature discovery in various domains. Identifying literature groundings and research gaps is an essential step during the ideation stage of the research lifecycle (Foster, 2004; Palmer et al., 2009). To assist users in efficiently navigating the vast amount of relevant literature, especially in unfamiliar domains, we introduced the design of Literature Node, as shown in Figure 3. Each of the Literature Nodes presents a specific research topic, with a collection of scholarly publications proposed by the preceding expert persona. The collection of papers is retrieved from an online publication database using the pipeline described in Section 3.3.2. When a user clicks on the node, a right-side panel is revealed with detailed information about each publication. The user will be able to delete irrelevant papers or add additional papers using a search engine-like interface.
@@
[…]
Three components: LiteratureNode ‘NLP Algorithms for Article Classification’, LiteratureNode ‘Text Classification Techniques in Machine Learning’, and a Literature Information side panel for the selected paper ‘Text Classification Algorithms: A Survey’. Upon selecting a paper from the LiteratureNode ‘Text Classification Techniques in Machine Learning’, the Literature Information side panel displays detailed information about the paper, including authors, publication year, venue, citation count, and an abstract summarizing the paper’s content and key findings.
@@
Critique Node [D2]: feedback from the perspective of experts. Based on a Persona Node and Literature Node, the user will be able to generate a follow-up Critique Node, as shown in component 4 of Figure 1. The Critique Node contains a user-editable critique of the initial research idea based on the given Literature Node and impersonation of the expert persona from the Persona Node. The critique nodes are further used to illustrate the process of how peer feedback is given from the perspective of a persona other than the researcher using the system. We use the design of the three types of nodes to model the process of peer discussions and knowledge solicitation with researchers or experts from different disciplines and domains.
@@
RQ Node and Research Outline Panel [D1]: iterative research ideation and refinement through expert feedback. Figure 4 demonstrates the RQ Node and Research Outline Panel feature. This feature allows users to iteratively improve their research ideas by generating and refining research questions (RQs) and their corresponding outlines. The design decision of making the system RQ-driven is due to the crucial role of RQ in research development and its nature of being iteratively refined during research ideation (Elio et al., 2011; Foster, 2004). When a user selects an RQ Node, the system automatically generates a Research Outline panel, which provides a structured overview of a potential research project derived from the RQ. The Research Outline Panel contains five sections: research question, literature review, research scenario, potential research outline, and expected outcomes. The literature review section provides a summary of relevant literature, prompting the user to think critically about the existing research landscape. The system then encourages the user to specify a detailed research scenario, with several system-generated suggestions. Based on this input, the system dynamically generates a comprehensive research outline with key sections such as motivation, objectives, methodology, expected outcomes, and potential limitations. This interactive process allows users to iteratively refine their research outline by reviewing and incorporating system-generated suggestions and insights.
3.2.2. Example Use Case Scenario — a System Walkthrough
Imagine Linh, an HCI researcher in Health Informatics, who became interested in exploring the potential application of LLMs in their familiar domain. Linh began by creating a new Research Question (RQ) Node in the mind map canvas to guide the exploration of “fine-tuning Large Language Models to promote preventive health behavior.” After writing down the idea in the initial RQ node, Linh clicked the action button on the node to generate follow-up nodes. The system then generated three Persona Nodes based on the initial idea, and Linh selected “Behavior Psychology Researcher,” edited its profile, and proceeded to generate more subsequent nodes. Three Literature Nodes were revealed, each containing a preview of suggested papers. Interested in “Promoting Health Behavior Change With AI,” Linh expanded that node, reviewed its papers, and edited the list. Next, Linh generated three Critique Nodes providing feedback on the initial idea; one focused on “Theoretical Framework” and referenced psychology concepts like the “Transtheoretical Model of Change (TTM).” By requesting more follow-up nodes, Linh discovered three revised versions of the initial RQ. Selecting one incorporating TTM, Linh opened a Research Outline Panel summarizing past work and prompting a detailed research scenario. After specifying “Fine-tuning LLMs to deliver personalized stage-targeted health behavior change messages,” the system generated a research outline table with sections such as motivation, objectives, methodology, and limitations, which Linh further refined to finalize the study plan.
@@ 3.2.2. Example Use Case Scenario — a System Walkthrough
[…]
Three components: Research Question Node A, Research Question Node B and a Research Outline Panel. The Research Question Node A presents a predefined question on ‘How articles can be classified into different categories using various classification algorithms and which evaluation metrics (accuracy, precision, recall, F1-score) provide the most comprehensive performance assessment ?’. The Research Question Node B presents ‘What are the most effective classification algorithms for categorizing articles, and how evaluation metrics such as accuracy, precision, recall and F1-score vary in assessing their performance?’, and allows the user to edit the research question. Research Question Node B has a red heart clicked on top of the node. Upon selecting the Research Question Node B, the Research Outline Panel on the right displays sections including ‘Research Question’, ‘Literature Review’, ‘Research Scenario’, ‘Potential Research Outline’, and ‘Expected Outcomes’, enabling users to structure and export their study.
@@ 3.3. PersonaFlow System Backend and Implementation
In this section, we detail the implementation details of the PersonaFlow  system backend [^2] https://github.com/yiren-liu/personaflow/. The backend uses GPT-4o for text generation, and is designed to provide on-demand results by generating and synthesizing expert personas, identifying relevant literature based on research questions and personas, providing critiques by impersonating the personas, and generating research outlines. The PersonaFlow system is developed as a web application in Typescript, utilizing ReactJS and TailwindCSS for the frontend. The interactive flow editor is implemented using React Flow [^3] https://github.com/wbkd/react-flow/. For the application backend, Python with FastAPI [^4] https://github.com/tiangolo/fastapi/ is employed as the RESTful API server framework. We present our detailed prompts used in Section A.5.
3.3.1. Synthesizing expert profiles
The system generates different personas to provide relevant literature and critique, based on the research context. The expert profiles can be generated under 3 different scenarios: 1) Initial User RQ: When an RQ is initially provided by the user, the system uses the RQ to prompt our LLM to generate three relevant personas that simulate real-world individuals to augment ideation and brainstorming; 2) Literature Node-based Persona: When a researcher is recommended a literature node relevant to their RQ, they can generate personas using the literature node to ensure the generation of relevant personas. The persona is generated by utilizing the abstracts from the previous literature, which are processed using a scientific sentence classification pipeline built with a fine-tuned BERT model (Cohan et al., 2019; Beltagy et al., 2019). This pipeline extracts the background, methodology, and conclusion of the research. Finally, it summarizes the literature node to provide context for generating personas; 3) Author Profiling-based Persona Generation: To leverage the expertise of scholarly authors whose papers were included in the literature node, we developed an author profiling-based persona generation. Using the Semantic Scholar API (Kinney et al., 2023), we extracted papers from authors whose work was being recommended in the literature node. We identified the top three papers for each first author (since first authors are typically the primary contributors and experts in the research area the paper addresses), based on cosine similarity555with a threshold score of .65 using FAISS (Douze et al., 2024). These papers were summarized and used as context to generate personas aiming to capture the expertise and background of the authors.
@@ 3.3.1. Synthesizing expert profiles
3.3.2. Expert perspective-driven Paper Retrieval
We designed and implemented an expert perspective-driven scholarly paper retrieval module to assist users in identifying the relevant papers with higher coverage. However, using a research question directly as a search query to retrieve papers from online databases often leads to very few or irrelevant results. As shown in Figure 5, the PersonaFlow system conducts two steps when retrieving relevant papers: 1) query decomposition to yield high coverage of papers under different sub-topics; 2) semantic-based re-ranking to reduce the number of papers and only present the topic-relevant ones to the user based on user’s prior research question. During the query decomposition step, we prompt the LLM to first break down the user’s initial research question into primary queries that capture various aspects of the main topic. Then, each of the primary queries will be decomposed into several secondary sub-queries to optimize the search results using the Semantic Scholar search API (Kinney et al., 2023). A similar approach has also been found to be used in existing work for optimizing literature discovery results (Zheng et al., 2024). For instance, the research question “How to address the lack of engagement of users using online art platforms by simulating different AI personas?” can be broken down into primary queries including “engagement in online art appreciation” and “multi-persona simulation using AI.” Subsequently, each primary query will be further decomposed into secondary queries that are shorter keywords. For example, the primary query “engagement in online art appreciation” will be decomposed into “user engagement,” “art appreciation” and “AI persona.” After the search results for each primary and secondary query are gathered, they are aggregated and re-ranked using a dense retrieval model (Chen et al., 2024) that aligns the retrieved papers with the user’s original research question. We take the top-k most relevant papers based on semantic similarity and present them to the user. During the later user studies, we present the top 10 most relevant papers to the user.
Figure 5. The method used for retrieving papers using agents with expert personas.
[…]
Figure illustrates an iterative workflow in which expert‑persona agents help retrieve research papers. The process begins with a box labeled “Research Question,” exemplified by the inquiry: “How to address the lack of engagement of users using online art platforms by simulating multi‑persona?” An arrow guides this question to a “Query Decomposition” panel where a thinking robot icon generates two primary queries—one about engagement in online art appreciation and another about multi‑persona simulation using AI. Each primary query is then dissected into more focused secondary queries, such as “user engagement,” “art appreciation,” and “AI persona.” These refined queries flow downward to the Semantic Scholar API, represented by a search‑bar icon, which returns stacks of papers. The retrieved papers pass into a re‑ranking module—visualized as a bar of colored circles—that filters and selects the top‑K papers for each topic. A looping arrow from this output back to the original research‑question box signals that the entire process can repeat, continually refining results with the aid of expert personas.
@@ 4. User Studies Evaluating Co-Creation with PersonaFlow
  1. RQ1: How do multiple perspectives generated by PersonaFlow affect users’ perception of the research ideation process and its outcomes?
  2. RQ2: How does the design of PersonaFlow support users’ critical thinking activities in research ideation?
  3. RQ3: How does the use of different PersonaFlow features impact users’ perceived experience and outcomes?
@@ 4.1. User Study Design
We conducted a two-phase exploratory user study to evaluate the PersonaFlow system. All studies were completed remotely online over Zoom, where participants were asked to share their screens. Study procedures were approved by the IRB of the researchers’ institution. Participants were compensated at \$20 per hour for all user studies.
@@
Phase 1: Task-based exploration with single and multiple expert perspectives. We designed a task-based exploratory user study to understand the impact of LLM-simulated experts on users’ research ideation process and how they interact with these personas. Simulating the scenario where the user seeks help from domain experts when exploring a research topic, we asked participants to develop their own topics of interest and utilize our systems for ideation support. We then analyze how the ideation outcome and user perception change temporally throughout the exploration session as participants interact with different expert perspectives. In order to examine our system’s utility in supporting interdisciplinary research ideation, we recruited 21 participants from 8 different research institutions and different backgrounds (e.g., HCI, Bio-informatics, Machine Learning, etc.) with prior experience in conducting interdisciplinary research. Detailed information about participants’ backgrounds can be found in Section A.1.
@@
Specifically, we focus on examining users’ interactions with multiple perspectives provided by different AI experts temporally throughout the ideation process. At the start of each session, the user was first asked to interact with a single AI expert and complete a single iteration of research ideation. We instructed the user to pick one AI persona out of the first batch of LLM synthesized personas, gather literature, and receive critiques until the first set of revised research questions is reached. The user was then provided with a post-survey form for quantitative evaluation of their experience and perceptions pertaining to the first iteration. Subsequently, the user was asked to interact with additional experts by building on the outcomes from the previous iteration and continue exploring by generating and involving multiple AI personas until they reached the second set of research questions. A post-survey was also provided inquiring about their experience and perceptions of the second iteration. By the end of the study, we conducted an exit interview to gather qualitative feedback on their experience and rationalization of their interactions when using the system. We chose not to run a controlled experiment with a single agent (e.g., ChatGPT) setup as baseline for two reasons: 1) Participants suggested during the formative study (Section 3.1), off-the-shelf LLM chatbots like ChatGPT tend to generate generic responses and are hard to prompt; 2) during the pilot of the study design, we found that each session would take more than 1.5 hour to complete, and including an additional baseline would drastically increase the cognitive burden and cause potential fatigue for the recruited participants. More importantly, our study aims mainly to understand users’ engagement and behavior changes during the ideation process, thus we decided to adopt a single session exploratory study which focuses on gather more in-depth insights into users’ interaction with multiple expert perspective.
@@
Phase 2: Optional open-ended free exploration. At the end of the exploratory study, we invited participants to an optional 30-minute free exploration session where they could freely use the system’s features with minimal guidance and no procedural constraints. Participants were free to build on their existing flow from the previous session or create new flows using another topic. In total, 10 participants enrolled in the open-ended exploratory study. This study session aims to provide a deeper understanding of how users would use the system in a real-world exploration setting, as well as the benefits and shortcomings of the PersonaFlow system.